by substitution:
1. y=3x
y-x=30
2. x=y
2x+2y=120

Please help meeee. :)

Answer by jim_thompson5910(35256)

y-x=30

3x-x=30 ... plug in y = 3x (ie replace y with 3x)

2x=30

x=30/2

x=15

----
y = 3x

y = 3(15)

y = 45

So x = 15 and y = 45

=======================================================
# 2

2x + 2y =120

2y + 2y =120 ... plug in x = y (ie replace x with y)

4y =120

y =120/4

y =30

So x = 30 and y = 30